# SplatVoxel: History-Aware Novel View Streaming without Temporal Training

### v1

@@ 4.1.2 Splat-to-Voxel Transfer
Given voxel grid $V$, we deposit each Gaussian primitive $\mathbf{G}_{k}$ to the set of $\{\mathbf{M}_{k}\}$ of nearest voxels based on the position $\mu_{k}$ using a distance kernel $\mathcal{K}$. That is, for voxel $V_{j}$, its weight associated with $\mathbf{G}_{k}$ is:
@@
where $\mathbf{x}_{j}$ represents the center of the $j$-th voxel (details in supplementary). We choose the adjacent 8 voxels for each splat in our implementaion. We note that the same splat depositing procedure can be used for Gaussian primitives originating from multi-view images, or warped from previous frames as will be detailed in Sec. 4.2. This seamlessly integrates all incoming splats which will be used to generate refined primitives with the downstream voxel transformer.
@@
$\mathbf{E}_{j}=\frac{1}{\mathbf{W}_{j}}\sum_{k\in\{\mathbf{G}_{k}\}}w_{jk}\alpha_{k}\mathbf{F}_{k},\quad\mathbf{W}_{j}=\sum_{k\in\{\mathbf{G}_{k}\}}w_{jk}\alpha_{k}$ (4)
where $w_{jk}$ is interpolation weight computed in Eqn. 3. We use the notation $k\in\{\mathbf{G}_{k}\}$ to denote an index into the set of all primitives, with associated weight, opacity, and feature.
@@
Once the splat features are deposited onto the voxels, we form the coarse-to-fine voxel hierarchy. Using the previously computed voxel features $\mathbf{E}_{j}$ as the fine voxel features, we now compute coarse voxel features $\mathbf{E}^{c}_{i}$. Let $\mathbf{N}_{i}$ be the set of fine voxels associated $i$-th coarse voxel. We derive coarse voxel features using a shallow multi-layer perceptron (MLP) and set coarse voxel weights as the sum of the associated fine voxel weights:
@@
We leverage geometric cues from coarse voxel weights $\mathbf{W}_{i}$, which encapsulate the spatial distribution and opacity of Gaussian primitives within each voxel, to efficiently cull empty and insignificant regions of the scene. We sort the coarse-level voxels by their voxel weights and retain only the top 20% with the largest weights as a sparse voxel representation of the scene. This sparsification is crucial for the downstream voxel transformer, and reduces the input token count to fewer than $10K$, which allows the voxel transformer to run in $15$ ms.
4.1.3 Sparse Voxel Transformer
From coarse voxel features, we treat each voxel feature as a token and reshape to 1D vectors. We train a transformer to process the set of coarse voxel features $\mathbf{E}^{c}$:
@@ 4.1.3 Sparse Voxel Transformer
The processed latent features $\{\mathbf{O}^{c}\}$ are then replicated from the coarse voxel grids within their corresponding fine voxel grids, producing latent features $\{\mathbf{O}\}$ at the fine voxel resolution. A shallow MLP subsequently generates refined Gaussian primitives by integrating the initial fine-level voxel attributes $\mathbf{E}_{j}$ with the fine-level transformer latent $\mathbf{O}_{j}$.
@@
4.1.4 Training
We train the SplatVoxel feed-forward networks on multi-view static scene datasets [82, 37] using a photometric loss combining Mean Squared Error (MSE) and perceptual LPIPS [81]:
@@ 4.1.4 Training
where $\hat{I}$ denotes ground-truth target image and $I$ the corresponding rendered target image from the predicted Gaussian primitives, which come from either the multi-view transformer or voxel transformer. We first train the multi-view transformer, then jointly both multi-view and voxel transformers together. We set $\lambda$ to 0.5 for the multi-view transformer and 4.0 for the voxel transformer. Training details are provided in the supplementary.
Table 2: Re10K Reconstruction Metrics. We train our model on the RealEstate10K dataset and compare it with feed-forward Gaussian Splatting methods. Baseline metrics are obtained from their offical papers.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
pixelSplat [5] 26.09 0.863 0.136
MVSplat [6] 26.39 0.869 0.128
GS-LRM [79] 28.10 0.892 0.114
@@
Figure 3: Static scene reconstruction. Qualitative comparison of our method with recent feed forward GS-LRM [79] on RealEstate10k [82]. Our method preserves high fidelity reconstructions and sharp details.
4.2 History-aware Streaming Reconstruction
In this section, we describe how our trained feed-forward SplatVoxel transformer enables zero-shot history-aware novel view streaming, without training on dynamic scenes.
Table 3: Novel View Streaming Metrics. We compare reconstruction quality of our model to prior methods. Our model differs from these baselines by being history-aware, in contrast to per-frame or temporal methods, and feed-forward rather than optimization-based. Our model outperforms baselines in reconstruction quality and ranks second in temporal smoothness (flicker), whereas the best-ranking baseline in flicker suffers a steep drop in image quality. We report the average running time of processing a single frame on Neural3DVideo at the resolution of $320\times 240$. GSLRM and Ours are tested on a single H100 card. 3DGS, 3DGStream, and 4DGS are tested on a single A100 GPU due to limitations with the H100 card. While the efficiency of these optimization-based methods can be enhanced, they still do not achieve a frame rate of 1 FPS. The best and second-best results are highlighted for clarity.
Category Method Time (s) 4-view LongVolumetricVideo [75] 2-view Neural3DVideo [26]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Flicker $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Flicker $\downarrow$
Per-frame Optimization 3DGS [22] 9.8 18.588 0.6368 0.4334 61.686 21.02 0.6234 0.4989 43.51
3DGStream [62] 3.8 16.970 0.4405 0.5040 12.436 14.24 0.3542 0.6016 6.554
@@ 4.2 History-aware Streaming Reconstruction
History-aware Feed Forward Ours 0.07 (non-keyframe) 25.564 0.8645 0.1242 7.782 27.41 0.8863 0.1040 2.916
@@
Figure 4: Streaming Reconstruction. We compare against per-frame methods (3DGS and GS-LRM) and temporal methods (4DGS and 3DGStream) on LongVolumetricVideo [75] with 4 input views and Neural3DVideo [26] with 2 input views. 3DGS and 3DGStream struggle with sparse input views, while 4DGS is susceptible to occlusion artifacts. GS-LRM is not designed to handle temporal sequences and prone to flicker, while our method mitigates flickering and produces realistic geometry especially in challenging occluded areas.
4.2.1 3D Warping
Due to scene motion, directly using Gaussian primitives from past frames results in lagging artifacts and fails to capture newly appearing objects. Therefore, we use 2D motion cues from pre-trained models to warp 3D Gaussian primitives from past frames to the current frame.
Triangulation
Given a Gaussian primitive with an initial position $\mu_{t^{\prime}}$ at a previous frame $t^{\prime}$, we estimate its corresponding 3D position $\mu_{t}$ at the current frame $t$ by performing triangulation on the 2D correspondance obtained by a pretrained 2D point tracking model [10]. In detail, we first compute its 2D projection $\{\mathbf{p}_{t^{\prime}}^{i}\mid i=1...N\}$ across $N$ input views. We then operate 2D point tracking on the projection points to obtain its corresponding 2D pixel position $\mathbf{p}_{t}^{i}$ in the current frame. The 3D position is subsequently recovered by finding the closest point to the $N$ camera rays originating from the 2D projections $\mathbf{p}_{t}^{i}$, which can be formulated as the following least-squares optimization problem:
where $\mathbf{o}_{i}$ and $\mathbf{d}_{i}$ denote the origin and direction of the $i$-th camera ray, respectively, and $\lambda_{i}$ is a scalar representing the depth along the ray. This least-squares problems can be further converted into solving an linear equation as
Embedded Deformation Graph
While the algorithm above can estimate 3D motion for any splat, doing so for all primitives is computationally expensive due to 2D tracking, and lacks robustness when splats are occluded. Instead, we use an embedded deformation graph [60, 42, 24] to introduce a spatial continuity prior during online motion estimation. We begin by selecting a subset of anchor splats that are consistently visible across multiple views, using farthest sampling [46]. We then construct the embedded deformation graph using a K-nearest neighbors (KNN) approach, binding each splat to its $K$-nearest splats and computing blend weights based on distance. To warp splats from the previous frame, we computed embedded deformations on the anchor splats using triangulation, and then propagate motion to all primitives through Linear Blend Skinning. For efficiency, we only maintain primitives from a set of past keyframes, and the contribution of these primitives are smoothly adjusted as older keyframes are replaced by newer ones to ensure seamless temporal transitions. Due to time constraints from querying 2D tracking model, in practice we use two keyframes placed 5 frames apart, each lasting for 10 frames.
4.2.2 Error-aware Temporal Fusion
During online reconstruction, we incorporate both historical Gaussian primitives warped from keyframes and new Gaussian primitives predicted from multi-view observation at the current frame. All of the collected Gaussian primitives are deposited to voxels (Sec. 4.1.2) and then processed by the voxel transformer (Sec. 4.1.3), producing refined, history-aware Gaussian primitives.
To further mitigate artifacts resulting from warping inaccuracies of historical primitives (e.g. warping error, appearance variattion, or new objects), we use an adaptive fusion weighting mechanism during splat fusion based on the per-pixel error between the past and current frame splats when rendered to the input views. In more detail, each historical splat is projected onto the input views to assess whether it produces a higher rendering error and appears in front of the rendered depth of the current frame splats. If a historical splat results in increased error in more than half of the input views, we set its fusion weight to zero to suppress unreliable contributions. Concurrently, current frame splats predicted from the multi-view transformer that increase rendering error are assigned zero weight. To enhance temporal continuity, we retain 50% of the voxel attributes from the past keyframe in static areas.
5 Experiments
In this section, we present evaluations of novel view synthesis on both static and dynamic scenes, along with ablation studies on framework design.
5.1 Feed-forward Novel View Synthesis
We first compare our proposed two-stage feed-forward networks with other feed-forward reconstruction methods on the RealEstate10k [82], following the trainining and testing setup in [5]. Since GS-LRM is not open-source, we re-implement it and report its reproduced performance. Table 2 presents a quantitative comparison based on commonly used novel view synthesis metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [70], and the perceptual metric LPIPS [80]. As shown in Fig. 3, our method captures finer details, such as thin structures (e.g., chairs) and reflections, outperforming the state-of-the-art feed-forward Gaussian splatting method, GS-LRM [79].
Additionally, we conduct quantitative comparison with GS-LRM on the DL3DV dataset [37] (Table 4). Our method outperforms GS-LRM by approximately 2dB in PSNR while maintaining a comparable transformer capacity with similar inference time, and using the same training setup. Qualitative comparisons and training details are provided in the Supplementary Document.
Table 4: DL3DV Reconstruction Metrics. We compare our approach with the state-of-the-art feed-forward Gaussian splatting method GS-LRM [37] and present ablation studies on our model designs.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time (ms)
GS-LRM 28.59 0.925 0.063 52.8
@@ 5.1 Feed-forward Novel View Synthesis
Ours (w/o Coarse-to-fine) 20.62 0.587 0.247 48.6
Ours (w/o Sparse Voxel) 29.69 0.926 0.060 72.0
Ours (3D CNN Downsample) 29.44 0.922 0.061 82.1
Ours (w/o Splat Feature) 29.40 0.924 0.061 49.2

### v2

@@ 4.1 Splat-to-Voxel Transfer
Given a dense high-resolution voxel grid $V$, we deposit each Gaussian primitive $\mathbf{G}_{k}$ to the set of $\{\mathbf{M}_{k}\}$ of nearest voxels based on the position $\mu_{k}$ using a distance kernel $\mathcal{K}$. That is, for voxel $V_{j}$, its weight associated with $\mathbf{G}_{k}$ is:
@@
where $\mathbf{x}_{j}$ represents the center of the $j$-th voxel. We choose the adjacent 8 voxels for each splat in our implementation.
@@
$\mathbf{E}_{j}=\frac{1}{\mathbf{W}_{j}}\sum_{k\in\{\mathbf{M}_{k}\}}w_{jk}\alpha_{k}[\mathbf{F}_{k};\mathbf{G}_{k}],\quad\mathbf{W}_{j}=\sum_{k\in\{\mathbf{M}_{k}\}}w_{jk}\alpha_{k}$ (5)
where $\mathbf{F}_{k};\mathbf{G}_{k}]$ denotes the concatenation of the feature vector $\mathbf{F}_{k}$ and Gaussian attributes $\mathbf{G}_{k}$. The interpolation weight $w_{jk}$ is computed as described in Eqn. 4. We use the notation $k\in{\mathbf{M}_{k}}$ to indicate that $k$ indexes over the set of all Gaussian primitives, each associated with a weight, Gaussian attributes, and a feature vector.
@@
Once the splat features are deposited onto the voxels, we construct a coarse-to-fine voxel hierarchy. Specifically, the high-resolution voxels are downsampled by a factor of $[d,h,w]$ to generate a low-resolution voxel grid that serves as the coarse scene representation. Sharing the same spatial topology, each coarse-level voxel directly corresponds to a $[d,h,w]$ block of fine-level voxels and can be restored by subdivision. We denote $\mathbf{N}_{i}$ as the set of $d\times h\times w$ fine voxels associated with the $i$-th coarse voxel. Using the previously computed voxel features $\mathbf{E}_{j}$ as the fine voxel features, we now compute coarse voxel features $\mathbf{E}^{c}_{i}$.
We derive coarse voxel features using a shallow multi-layer perceptron (MLP) on the concatenation of the fine features and set coarse voxel weights as the sum of the associated fine voxel weights:
@@
Voxel Sparsification
We leverage geometric cues from coarse voxel weights $\mathbf{W}_{i}$, which encapsulate the spatial distribution and opacity of Gaussian primitives within each voxel, to efficiently cull empty and insignificant regions of the scene. We sort the coarse-level voxels by their voxel weights and retain only the top 20% with the largest weights as a sparse voxel representation of the scene. This sparsification is crucial for the downstream voxel transformer, and reduces the input token count to fewer than $10K$, which allows the voxel transformer to run in $15$ ms. Note that the high-resolution fine-level voxels are also sparsified accordingly if their corresponding coarse-level voxels are culled.
4.2 Sparse Voxel Transformer
From coarse voxel features, we treat each voxel feature as a token and reshape them to 1D vectors. We train a transformer to process the set of coarse voxel features $\mathbf{E}^{c}$:
@@ 4.2 Sparse Voxel Transformer
The processed latent features $\{\mathbf{O}^{c}\}$ are then replicated from the coarse voxel grids within their corresponding fine voxel grids, producing latent features $\{\mathbf{O}_{j}\}$ at the fine voxel resolution. A shallow MLP subsequently generates refined Gaussian primitives by integrating the initial fine-level voxel attributes $\mathbf{E}_{j}$ with the fine-level transformer latents $\mathbf{O}_{j}$.
@@
4.3 Training
We train the SplatVoxel feed-forward networks on multi-view static scene datasets zhou2018rel10k; ling2024dl3dv using a photometric loss combining Mean Squared Error (MSE) and perceptual LPIPS (zhang2018perceptual):
@@ 4.3 Training
where $\hat{I}$ denotes ground-truth target image and $I$ the corresponding rendered target image from the predicted Gaussian primitives, which come from either the multi-view transformer or voxel transformer. We set $\lambda$ to 0.5 for the multi-view transformer and 4.0 for the voxel transformer. We train our full-scale model for cross-dataset generalization on the DL3DV dataset ling2024dl3dv with a batch size of 128 for a total of 300K iterations using a two-stage training strategy. In the first stage, we train the multi-view Transformer backbone for 200K iterations, followed by joint fine-tuning of both the multi-view and voxel Transformers for an additional 100K iterations. Network architecture and further training details are provided in the supplementary material (Sec. A.1 and Sec. B).
Table 1: Quantitative Comparisons on the RealEstate10K Dataset. Our method achieves state-of-the-art performance among recent feed-forward 3D Gaussian Splatting methods.
[…]
pixelSplat charatan2024pixelsplat 26.09 0.863 0.136
MVSplat chen2024mvsplat 26.39 0.869 0.128
TranSplat zhang2025transplat 26.69 0.875 0.125
HiSplat tang2024hisplat 27.21 0.881 0.117
OmniScene wei2025omni 26.19 0.865 0.131
DepthSplat xu2025depthsplat 27.47 0.889 0.114
GS-LRM zhang2025gslrm 28.10 0.892 0.114
@@
Table 2: Quantitative Comparisons on the DL3DV Dataset. We build our Voxel Transformer on GS-LRM ling2024dl3dv and present ablation studies on model design. For a fair comparison, we use 12 multi-view Transformer layers with our Voxel Transformer to match the size and runtime of GS-LRM, which uses 24 layers.
@@
GS-LRM 28.59 0.925 0.063 52.8
+ Non-learning Fusion 12.57 0.357 0.741 33.8
+ Ours (w/o Coarse-to-fine) 20.62 0.587 0.247 48.6
+ Ours (w/o Sparse Voxel) 29.69 0.926 0.060 72.0
+ Ours (w/ 3D CNN) 29.44 0.922 0.061 82.1
+ Ours (w/o Splat Feature) 29.40 0.924 0.061 49.2
+ Ours 30.61 0.935 0.052 52.5
5 Zero-shot Streaming Fusion
In this section, we demonstrate an application of the Sparse Voxel Transformer, trained solely on static scenes, to enable zero-shot history-aware novel view streaming without requiring any training on dynamic scenes. Please refer to the Appendix (Sec. A.3) for full implementation details.
3D Warping
Given a Gaussian primitive with an initial position $\mu_{t^{\prime}}$ at a previous frame $t^{\prime}$, we first estimate its corresponding 3D position $\mu_{t}$ at the current frame $t$ by performing triangulation on the 2D correspondence obtained by a pre-trained 2D point tracking model doersch2023tapir. In detail, we first compute its 2D projection $\{\mathbf{p}_{t^{\prime}}^{i}\mid i=1...N\}$ across $N$ input views. We then run 2D point tracking on the projected points to obtain their corresponding 2D pixel position $\mathbf{p}_{t}^{i}$ in the current frame. The 3D position is subsequently recovered by finding the closest point to the $N$ camera rays originating from the 2D projections $\mathbf{p}_{t}^{i}$. We adopt embedded deformation graph sumner2007ed_graph; newcombe2015dynamicfusion; lei2024mosca to propagate reliable motion estimates from a sparse set of anchor points to the entire scene. Anchor points are selected via farthest point sampling qi2017pointnet++, and an embedded deformation graph is constructed by connecting each splat to its $K$ nearest anchor points, with blend weights computed based on spatial distances. For efficiency, we only maintain primitives from a set of past keyframes, and the contribution of these primitives are smoothly adjusted as older keyframes are replaced by newer ones to ensure seamless temporal transitions.
Error-aware Fusion
During streaming reconstruction, we incorporate both historical Gaussian primitives warped from keyframes and new Gaussian primitives predicted from multi-view observations at the current frame. All of the collected Gaussian primitives are deposited to voxels (Sec. 4.1) and then processed by the voxel transformer (Sec. 4.2), producing refined, history-aware Gaussian primitives. To mitigate artifacts resulting from warping inaccuracies of historical primitives (e.g. warping error, appearance variation, or new objects), we use an adaptive fusion weighting mechanism during splat fusion based on the per-pixel error between the past and current frame splats when rendered to the input views.
Table 3: Quantitative Comparison Under Varying Input Views on the DL3DV Dataset. Trained with 4 input views, our method consistently outperforms GS-LRM across different input-view settings at inference.
Method 2 Views 4 Views 8 Views 16 Views
@@ 5 Zero-shot Streaming Fusion
GS-LRM 18.99 0.810 0.135 28.59 0.925 0.063 22.51 0.879 0.106 20.12 0.792 0.177
Ours 26.32 0.876 0.093 30.34 0.934 0.054 28.98 0.930 0.060 26.06 0.889 0.093
Figure 2: Static Scene Reconstruction. Qualitative comparison of our method with GS-LRM zhang2025gslrm on RealEstate10k zhou2018rel10k. Our method preserves high fidelity reconstructions and sharp details.
6 Experiments
In this section, we present evaluations of feed-forward novel view synthesis on both static and dynamic scenes, along with ablation studies on our splat-voxel representation designs.
6.1 Feed-forward Novel View Synthesis
We benchmark our method on two widely used datasets, RealEstate10K zhou2018rel10k and DL3DV ling2024dl3dv, which cover both indoor scenes and unbounded large-scale environments.
Following the training and testing protocol of charatan2024pixelsplat, our method achieves state-of-the-art performance on RealEstate10K, as shown in Table 2. As illustrated in Fig. 2, our method reconstructs finer details such as thin structures (e.g., chairs) and reflections compared to the strongest baseline, GS-LRM zhang2025gslrm. Additional qualitative comparisons and training details are provided in the Appendix (Sec. B and Sec. D).
We then demonstrate the ability of our method to enhance existing feed-forward 3D Gaussian Splatting approaches through a comparison on the DL3DV dataset ling2024dl3dv with GS-LRM zhang2025gslrm. Table 2 demonstrates that our method outperforms GS-LRM by approximately 2 dB in PSNR, while maintaining a comparable number of network parameters, similar inference time, and the same training configuration. Our method also demonstrates strong generalization capability across varying input views. Trained with 4 input views on the DL3DV dataset, our method consistently outperforms GS-LRM across awide range of input configurations as shown in the Table 3. These results further highlight the strong generalization ability of our method for long-sequence, large-scale scene reconstruction. A compairson with Gaussian Graph Network zhang2024gaussianGNN is provided in the appendix Table 13, which state-of-the-art method for handling diverse input views.
Table 4: Novel View Streaming Metrics. We compare reconstruction quality of our model to prior methods on two multi-view video datasets with sparse-view input. We report the average running time of processing a single frame on Neural3DVideo at the resolution of $320\times 240$. The best and second-best results are highlighted for clarity.
Method Time (s) Neural3DVideo li2022neural3DV LongVolumetricVideo xu2024longvolcap
[…]
3DGS kerbl20233dgs 9.8 21.02 0.6234 0.4989 43.51 18.56 0.6368 0.4334 61.69
3DGStream sun20243dgstream 3.8 14.24 0.3542 0.6016 6.554 16.97 0.4405 0.5040 12.44
4DGS wu20244d 6.0 23.16 0.7812 0.2079 2.746 18.75 0.6361 0.3554 6.528
GS-LRM zhang2025gslrm 0.04 21.80 0.8488 0.1278 5.714 25.49 0.8558 0.1377 8.463
Ours 0.07 (non-keyframe) 27.41 0.8863 0.1040 2.916 25.56 0.8645 0.1242 7.782
@@ 6.1 Feed-forward Novel View Synthesis
Figure 3: Qualitative Comparison of Novel View Streaming. We compare against per-frame methods (3DGS and GS-LRM) and temporal methods (4DGS and 3DGStream) on LongVolumetricVideo xu2024longvolcap with 4 input views and Neural3DVideo li2022neural3DV with 2 input views.
Table 5: Comparison of GS-LRM with Simple 3D Warping and Fusion Strategy. Incorporating historical information for streaming reconstruction remains challenging due to accumulated errors, whereas our Fuse-and-Refine network learns to effectively mitigate this limitation.
@@
GS-LRM 21.80 0.8488 0.1278 5.714
GS-LRM + 3D Warping 21.40 0.8250 0.1419 2.163
GS-LRM + 3D Warping + Fusion (Concat + Dropout) 19.52 0.7429 0.2502 3.983
GS-LRM + 3D Warping + Ours (Fuse-and-Refine) 27.41 0.8863 0.1040 2.916
Figure 4: Temporal Slice Comparison. We extract a horizonal line in the ground truth and generated videos to visualize temporal stability. Applying GS-LRM independently to each frame results in noticeable flickering, whereas our method achieves greater temporal consistency and improved occlusion robustness through temporal fusion and refinement.
6.2 Novel View Streaming
We conduct a cross-dataset evaluation of our method on dynamic scene reconstruction, demonstrating its generalization capability in enhancing reconstruction quality and improving temporal coherence. We train both GS-LRM and our model on DL3DV, and compare our method with baselines on the novel view streaming task on two dynamic scene datasets, Neural3DVideo li2022neural3DV and LongVolumetricVideo xu2024longvolcap. We use 2 input views for Neural3DVideo and 4 views for LongVolumetricVideo. For quantitative evaluation of temporal coherence, we introduce a flicker metric that complements the per-frame novel view reconstruction scores. Temporal flicker is quantified by measuring feature-space variations over time. For predicted frames $I_{t}$ and ground truth $I_{t}^{*}$, we compute
$\text{Flicker}_{t}=\big|D_{t}-D_{t}^{*}\big|,$
where $D_{t}=\|\phi(I_{t})-\phi(I_{t-1})\|_{2}$ and $D_{t}^{*}=\|\phi(I_{t}^{*})-\phi(I_{t-1}^{*})\|_{2}$ denote temporal feature differences extracted from a pretrained VGG network simonyan2014vgg. The final flicker score is computed by averaging $\text{Flicker}_{t}$ over the entire sequence. Lower values indicate smoother temporal consistency.
As presented in Table 4, our method achieves the highest per-frame novel view synthesis performance across all three evaluation metrics on both datasets with the second-best flicker scores. In comparison to the state-of-the-art per-frame feed-forward method GS-LRM, our approach qualitatively reconstructs geometry and appearance with greater accuracy and finer details (Fig. 3) Our method reduces flickering through history-aware modeling, visualized with temporal slices (Fig. 4). Although 4D Gaussian Splatting yields slightly better flicker scores, it produces significantly poorer novel views and its costly temporal optimization is impractical for online processing of long sequences.
We further apply 2D tracking and 3D warping to GS-LRM to demonstrate that directly warping past reconstructions introduces significant errors. As shown in Table 5, GS-LRM + 3D Warping fails to handle new content and accumulates warping errors over time. A simple fusion baseline (Concat + Dropout) mitigates primitive growth by discarding 50% per frame but lacks refinement and often removes accurate primitives.
For the reported running time in Table 4, GS-LRM and ours are tested on a single H100 card. 3DGS, 3DGStream, and 4DGS are tested on a single A100 GPU due to limitations with the H100 card. While the efficiency of these optimization-based methods can be enhanced, they still do not achieve a frame rate of 1 FPS. We apply 2D tracking only to keyframes sampled every 5 frames, where each keyframe requires 0.35 s, while non-keyframes take 0.07 s. A detailed runtime breakdown of our method for both static and dynamic scenes is presented in Table 6.
Table 6: Inference Runtime Breakdown of Our Method for Static and Dynamic Reconstructions.
Static Scenes ($\times256$ Resolution)
Multiview Transformer Splat-to-Voxel Transfer Voxel Transformer Rendering
Time (s) 0.026 0.006 0.015 0.003